OpenStudy (anonymous):

Compute the derivative for f(x) = sin^2 xcos x

myininaya (myininaya):

f'(x)=(sin^2x)'cosx+(cosx)'sin^2x

OpenStudy (anonymous):

product rule rules!

myininaya (myininaya):

who needs quotient when you have the product rule

myininaya (myininaya):

you can always rewrite a quotient as a product

OpenStudy (anonymous):

I don't have that as one of my answers in my homework. Is there any more breaking down or extra step?

myininaya (myininaya):

' means you have to find the derivative

myininaya (myininaya):

i didnt do the whole problem for you

myininaya (myininaya):

i just gave u the first step

myininaya (myininaya):

what is (cosx)'=??

myininaya (myininaya):

what is (sin^2x)'=??

myininaya (myininaya):

fill in these blanks above

OpenStudy (dumbcow):

i think you missed an x f(x) = x*sin(x)^2*cos(x)

myininaya (myininaya):

I don't see an x up there?

myininaya (myininaya):

I see f(x)=(sinx)^2*cosx